# Mathematics with the knowledge predicate ${\mathcal{K}}$ satisfied by every currently known theorem Apoloniusz Tyszka

### v18

@@ Abstract
${\mathcal{K}}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known theorems. The set ${\mathcal{K}}$ is time-dependent, publicly available, and contains theorems both from formal and constructive mathematics. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}$. Mathematical statements with known constructive proofs exist in ${\mathcal{K}}$ separately and form the set ${\mathcal{K}}_{\textstyle c}\subseteq{\mathcal{K}}$. We assume that mathematical sets are atemporal entities. They exist formally in $ZFC$ theory although their properties can be time-dependent (when they depend on ${\mathcal{K}}$) or informal. Algorithms always terminate. We explain the distinction between algorithms whose existence is provable in $ZFC$ and constructively defined algorithms which are currently known. By using this distinction, we obtain non-trivially true statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ that belong to constructive and informal mathematics and refer to the current mathematical knowledge on ${\mathcal{X}}$. We present a non-trivially true statement in computability (which may be falsified) in which the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge appears exclusively inside the statement.
Key words and phrases: constructive mathematics, constructively defined algorithms, current mathematical knowledge, informal mathematics, known algorithms, time-dependent notion of truth in constructive mathematics.
@@ 8 Subsets of $\mathbb{N}$ and their threshold numbers
9 A theorem in the classical computability theory
@@
([18, p. 118]). There is a limit-computable function $\beta_{1}:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\delta_{1}:\mathbb{N}\to\mathbb{N}$.
@@
Flowchart 1.  A semi-algorithm which computes $\beta_{1}(n)$ in the limit.
@@
10 A statement in computability theory with the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge
${\mathcal{K}}_{f}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known formal theorems. The set ${\mathcal{K}}_{f}$ is time-dependent and publicly available. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}_{f}$. In the statement: It is unknown whether or not there are infinitely many twin primes, the predicate ${\mathcal{K}}_{f}$ appears exclusively outside a formal mathematical statement.
In the statement: It is unknown whether or not there are infinitely many twin primes, the predicate ${\mathcal{K}}_{f}$ appears exclusively outside a formal mathematical statement. Statement 11 is non-trivially true, contains the predicate ${\mathcal{K}}_{f}$, and will be false when someone proves Conjecture 2. Since the function $\beta_{1}$ in Theorem 2 is not computable, Statement 11 does not follow from Theorem 2. Ignoring the epistemic condition in Statement 11, Statement 11 follows from Theorem 2. Statement 11 does not follow from any widely known mathematical theorem. Statement 11 shows that there is a non-trivially true statement in computability (which may be falsified) in which the predicate ${\mathcal{K}}_{f}$ appears exclusively inside the statement.
@@
There is a limit-computable function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
@@
This is proved in [23]. The term "dominated" in the title of [23] means "eventually dominated". For every $n\in\mathbb{N}$, we define $\beta(n)$ as the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. Flowchart 2 describes a semi-algorithm which computes $\beta(n)$ in the limit.
Flowchart 2.  A semi-algorithm which computes $\beta(n)$ in the limit.
Corollary 3.
The proof of Statement 11 shows that Statement 11 follows from some mathematical theorem and the following conjunction: ((The function $\beta$ is computable) $\not\in{\mathcal{K}}_{f}$) $\wedge$ ((The function $\beta$ is uncomputable) $\not\in{\mathcal{K}}_{f}$).

### v19

@@ Abstract
${\mathcal{K}}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known theorems. The set ${\mathcal{K}}$ is time-dependent, publicly available, and contains theorems both from formal and constructive mathematics. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}$. Mathematical statements with known constructive proofs exist in ${\mathcal{K}}$ separately and form the set ${\mathcal{K}}_{\textstyle c}\subseteq{\mathcal{K}}$. We assume that mathematical sets are atemporal entities. They exist formally in $ZFC$ theory although their properties can be time-dependent (when they depend on ${\mathcal{K}}$) or informal. Algorithms always terminate. We explain the distinction between algorithms whose existence is provable in $ZFC$ and constructively defined algorithms which are currently known. By using this distinction, we obtain non-trivially true statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ that belong to constructive and informal mathematics and refer to the current mathematical knowledge on ${\mathcal{X}}$. We present non-trivially true statements in computability with the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge, which do not express the current and formal mathematical knowledge and may be falsified.
Key words and phrases: constructive mathematics, constructively defined algorithms, current mathematical knowledge, informal mathematics, known algorithms, time-dependent notion of truth in mathematics.
@@ 8 Subsets of $\mathbb{N}$ and their threshold numbers
9 Classical computability theory
@@
([18, p. 118]). There exists a limit-computable function $\beta_{1}:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\delta_{1}:\mathbb{N}\to\mathbb{N}$.
@@
Flowchart 1
A semi-algorithm which computes $\beta_{1}(n)$ in the limit
@@
For every $n\in\mathbb{N}$, we define $\beta(n)$ as the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
Theorem 3.
The function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
This is proved in [23]. The term "dominated" in the title of [23] means "eventually dominated". Flowchart 2 describes a semi-algorithm which computes $\beta(n)$ in the limit.
Flowchart 2
A semi-algorithm which computes $\beta(n)$ in the limit
10 Non-trivially true statements in computability with the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge, which do not express the current and formal mathematical knowledge and may be falsified
${\mathcal{K}}_{f}$ denotes both the knowledge predicate satisfied by every currently known formal theorem and the finite set of all currently known formal theorems. The set ${\mathcal{K}}_{f}$ is time-dependent and publicly available. Any formal theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}_{f}$. The statement $\neg{\mathcal{K}}_{f}(P\neq NP)\wedge\neg{\mathcal{K}}_{f}(P=NP)$ is true and does not justify the title of the section.
@@
There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
@@
It follows from Theorem 3 by taking $f=\beta$ and the following conjunction:
$(({\rm The~{}function~{}}\beta{\rm~{}is~{}computable)~{}}\not\in{~{}\mathcal{K}}_{f})~{}\wedge~{}(({\rm The~{}function~{}}\beta{\rm~{}is~{}uncomputable)~{}}\not\in~{}{\mathcal{K}}_{f})$
Statement 12.
For every computable function $g:\mathbb{N}\to\mathbb{N}$, there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ of unknown computability such that $f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation and $f(n)>g(n)$ for every $n\in\mathbb{N}$.
Proof.
It follows from Theorem 3 by taking $f=g+1+\beta$ and the following conjunction:
$(({\rm The~{}function~{}}\beta{\rm~{}is~{}computable)~{}}\not\in{~{}\mathcal{K}}_{f})~{}\wedge~{}(({\rm The~{}function~{}}\beta{\rm~{}is~{}uncomputable)~{}}\not\in~{}{\mathcal{K}}_{f})$
Both Statement 11 and Statement 12 justify the title of the section. Statements 11 and 12 are non-trivially true, contain the predicate ${\mathcal{K}}_{f}$, and will be false when someone proves Conjecture 2. Since the function $\beta_{1}$ in Theorem 2 is not computable, Statements 11 and 12 do not follow from Theorem 2. Ignoring the epistemic condition in Statements 11 and 12, they follow from Theorem 2 by taking $f=\beta_{1}$ or $f=g+1+\beta_{1}$.